# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

### v2

@@ C.1 Pre-training and Fine-tuning Hyperparameters
All experiments are implemented in PyTorch using HuggingFace Transformers. Pre-training uses the built-in Trainer with DeepSpeed ZeRO-2 on multi-GPU data parallelism; fine-tuning is run as single-GPU jobs (one job per GPU) using the same Trainer API (or strategy-specific subclasses).
Pre-training on Enamine675M (MultiScaler3375M).
We pre-train a 16-layer LLaMA-style decoder initialized from a local config (i.e., random initialization; no pretrained weights are loaded). The training data are pre-tokenized and stored as HuggingFace datasets on disk with train/validation splits. We use FP16 training with a cosine learning-rate schedule and warmup, and evaluate/save checkpoints periodically during training.
Fine-tuning (MoleculeNet and MoleculeACE).
We fine-tune the pretrained backbone with fingerprint injection (fp_dim=2048) and a lightweight supervised head. All downstream runs use FP16, evaluate every epoch, and select the best checkpoint by the primary validation metric (higher-is-better for AUROC; lower-is-better for RMSE). We also tested two dropout application schemes in the MoleculeACE fine-tuning pipeline: (1) head-only dropout (applied only in the supervised head), and (2) backbone-synced dropout (also overriding backbone attention_dropout/hidden_dropout and all nn.Dropout modules). Both schemes share the same dropout-rate grid in Table 7.
Ablation compute budget.
For ablations (e.g., single-scale tokenization or removing fingerprint injection), we keep the optimizer, LR schedule, batch size, and training length the same as the corresponding main setting, and only change the ablated component.
Table 5: Pre-training hyperparameters on Enamine675M (MultiScaler3375M).
Setting Value
@@
Backbone config 16 layers; hidden size 720; 8 heads; MLP 2880
Context length 4096 (max_position_embeddings)
Initialization new model from local config (no pretrained weights)
Tokenizer CamS-tokenizer; embedding resized to tokenizer.vocab_size
Training instances Multi-scale tokenized corpus (MultiScaler3375M; 5 views per molecule)
Parallelism torchrun 8 processes + DeepSpeed ZeRO stage 2
Precision FP16 (fp16=true in Trainer and DeepSpeed)
@@
Gradient accumulation 2
Global batch size 4096 (256 $\times$ 8 $\times$ 2)
Optimizer AdamW (Transformers Trainer default)
Learning rate $1\times 10^{-4}$
@@
LR schedule cosine; warmup_steps =5000=5000
Gradient clipping 1.0 (DeepSpeed gradient_clipping)
Training length 10 epochs in planning (early stop after 1/10 epoch in actrual excution)
Eval / save / log eval_steps =33000=33000; save_steps =33000=33000; logging_steps =3300=3300
Checkpoint cap save_total_limit =100=100
Table 6: Fixed fine-tuning setup for MoleculeNet and MoleculeACE.
@@
Task type classification / regression (per dataset) regression
Data split provided scaffold splits (same as KPGT), fold $\in\{0,1,2\}$ provided scaffold splits (same as KPGT)
Epochs 80 (search phase) 50
Batch size 32 per GPU (default) 32 per GPU (default; a small subset rerun with 64)
Grad accumulation 1 (default) 1(default)
Optimizer AdamW (Trainer default) AdamW (Trainer default)
LR schedule constant_with_warmup, warmup_ratio =0.1=0.1 constant_with_warmup, warmup_ratio =0.1=0.1
Evaluation per epoch per epoch
Primary metric AUROC / macro-AUROC; RMSE for regression tasks RMSE (overall_RMSE)
Table 7: Fine-tuning hyperparameter search space (shared).
Hyperparameter Candidates
Base grid (48 combos) learning_rate $\in\{10^{-6},3\!\times\!10^{-6},10^{-5},3\!\times\!10^{-5}\}$
dropout_rate $\in\{0,0.05,0.1,0.2\}$; weight_decay $\in\{0,10^{-6},10^{-4}\}$
Fine-tune strategy normal, l2sp, llrd, flag, reinit
LLRD layer_decay $\in\{0.95,0.9,0.85,0.8\}$
L2SP weight $\in\{10^{-1},10^{-2},10^{-3},10^{-4}\}$
FLAG $\epsilon\in\{10^{-3},3\!\times\!10^{-3},5\!\times\!10^{-3},10^{-2}\}$
Reinit reinit_num_layers $\in\{1,2,3,4\}$
Appendix D Detailed Experimental Results
@@
Appendix E Interpretability Details for Activity-Cliff Attention Analysis
Overview. This section corresponds to the activity-cliff attention analysis in Sec. 4.3. It supplements the main text with implementation-level details of (1) activity-cliff pair construction and differential/shared atom identification (Appendix E.1), Alg.7, (2) mapping atom-level diff/shared labels to CamS tokens within each scale region (Appendix E.2), and (3) attention extraction and metric computation/aggregation for MDTA/MSTA and Rel-DTAP (Appendix E.3).
E.1 Activity-cliff Pair Construction and Differential/Shared Atom Identification
@@
E.2 Mapping Atom-level Diff/Shared Labels to Tokens in Each Scale Region
@@
Let $\mathcal{A}_{\Delta}$ be the differential-atom set for this molecule from Appendix E.1. We assign a token-level differential indicator by an any-diff rule: a token at position $t$ is marked as differential iff $S_{t}\cap\mathcal{A}_{\Delta}\neq\varnothing$; otherwise it is marked as shared/non-differential. Thus a token covering multiple atoms is counted as differential if it contains at least one differential atom.
@@
E.3 Attention Extraction and Metric Computation
@@
Within a given scale region $s$ (defined by [CONCAT] boundaries), let $p_{i}$ denote the (renormalized) attention weight on token $i$ in that region, and let $d_{i}\in\{0,1\}$ be its differential indicator from Appendix E.2. We define the mean differential-token attention and mean shared-token attention as

### v3

@@ C.1 Pre-training and Fine-tuning Implementation
All experiments are implemented in PyTorch using HuggingFace Transformers. Pre-training utilizes the Trainer API with DeepSpeed ZeRO-Stage 2 for multi-GPU data parallelism. Fine-tuning runs as single-GPU jobs using strategy-specific trainer subclasses.
Pre-training Configuration.
We train a 16-layer LLaMA-style decoder initialized from scratch (random initialization). The training corpus (Enamine675M augmented to $\sim$3.4B views) is pre-tokenized and stored as memory-mapped datasets. We use FP16 mixed precision, a cosine learning rate schedule with warmup, and periodic evaluation. The specific hyperparameters are listed in Table 5.
Fine-tuning Protocol.
We fine-tune the pre-trained backbone using the Dual-Path Strategy (Section 3.2), injecting fingerprints (specifically RDKit topological fingerprints, default parameters, 2048 bits) via a linear projection layer (fp_dim=2048). We employ two dropout schemes during fine-tuning (searched via grid search): (1) Head-only Dropout: Applied only to the task-specific prediction head. (2) Backbone-synced Dropout: Applied to the head and overriding all backbone dropout rates (attention and hidden states). Metrics are task-dependent: AUROC for classification (higher is better) and RMSE for regression (lower is better).
Table 5: Pre-training hyperparameters on Enamine675M.
Hyperparameter Value
@@
Backbone Architecture LLaMA Decoder (16 layers, 720 hidden size, 8 heads, 2880 MLP dim)
Context Length 4096 tokens
Initialization Random (from local config)
Tokenizer CamS-Tokenizer (Vocab size $\approx$ 67K)
Data Augmentation 5 views per molecule (4 single-scale + 1 multi-scale)
Parallelism DeepSpeed ZeRO Stage 2 (8 GPUs)
Precision FP16 Mixed Precision
@@
Gradient Accumulation 2 steps
Global Batch Size 4096 sequences ($256\times 8\times 2$)
Optimizer AdamW ($\beta_{1}=0.9,\beta_{2}=0.999$)
Peak Learning Rate $1\times 10^{-4}$
@@
LR Schedule Cosine with 5000 warmup steps
Gradient Clipping 1.0
Training Duration Two-stage training: Stage 1 (1 epoch) + Stage 2 (0.1 epoch with adjusted LR). Total samples seen $\approx$ 3.7B (where 1 epoch = 3.4B augmented views).
Table 6: Fine-tuning setup for downstream tasks.
@@
Task Type Classification / Regression Regression
Split Strategy Scaffold Split (KPGT protocol) Scaffold Split (Benchmark default)
Epochs 80 50
Batch Size 32 (per GPU) 32 (per GPU)
Optimizer AdamW AdamW
LR Schedule Linear decay (warmup ratio 0.1) Linear decay (warmup ratio 0.1)
Metric ROC-AUC (Cls) / RMSE (Reg) RMSE
Table 7: Hyperparameter search space for fine-tuning.
Hyperparameter Search Grid
Learning Rate $\{1\!\times\!10^{-6},3\!\times\!10^{-6},1\!\times\!10^{-5},3\!\times\!10^{-5}\}$
Dropout Rate $\{0.0,0.05,0.1,0.2\}$
Weight Decay $\{0.0,1\!\times\!10^{-6},1\!\times\!10^{-4}\}$
Fine-tune Strategy Standard, L2SP, LLRD, FLAG, Reinit
C.2 Benchmark Task Descriptions
We evaluate CamS-LLaMA on two complementary benchmarks to assess both general property prediction and structural sensitivity.
MoleculeNet (General Properties) (Wu et al., 2018).
This benchmark covers a diverse range of molecular properties. We select 11 datasets that cover physiology, biophysics, and physical chemistry. Consistent with our results table, the tasks are:
  • Classification Tasks (8):
    • BACE: Inhibition of $\beta$-secretase (a key Alzheimer’s therapeutic target).
    • BBBP: Blood-brain barrier penetration (permeability).
    • ClinTox: Clinical toxicity (distinguishing FDA-approved drugs from toxic compounds).
    • Estrogen: Estrogen receptor ($\alpha$, $\beta$) binding activity (endocrine disruption potential).
    • Metstab: Metabolic stability (half-life duration in liver microsomes).
    • SIDER: Adverse drug reactions (side effects) of marketed medicines.
    • ToxCast: High-throughput toxicology screening data.
    • Tox21: Toxicity testing across 12 biological targets (nuclear receptors/stress pathways).
  • Regression Tasks (3):
    • ESOL: Water solubility (log solubility in mols per litre).
    • FreeSolv: Hydration free energy (experimental vs calculated).
    • Lipo: Lipophilicity (octanol/water distribution coefficient, logD).
All datasets follow the scaffold splitting strategy (as used in KPGT) to strictly evaluate the model’s generalization capability to chemically distinct structures.
MoleculeACE (Activity Cliffs) (Van Tilborg et al., 2022).
MoleculeACE is designed to stress-test models on Activity Cliffs—pairs of molecules with high structural similarity but large differences in potency. It consists of 30 datasets derived from ChEMBL, each targeting a specific biological protein target (e.g., CHEMBL204, CHEMBL240).
  • Challenge: Unlike MoleculeNet, which often rewards global scaffold recognition, MoleculeACE requires the model to identify fine-grained structural edits (e.g., methylation, halogenation) that trigger drastic activity shifts.
  • Metric: Performance is measured by RMSE on the test set (scaffold split). Lower RMSE indicates better ability to capture the non-smooth structure-activity landscape.
Evaluation Protocols and Statistical Reporting.
[…]
Appendix D Comparison with Large-scale SMILES FMs
[…]
Table 8: Comparison with a SOTA SMILES-based Natural Language Science Foundation Models (NatureLM (Xia et al., 2025)). Results for NatureLM are taken directly from the original paper (Table 19). CamS-LLaMA ($\sim$0.1B) significantly outperforms NatureLM (up to 56B parameters) on common benchmarks, demonstrating superior representational efficiency.
Model Params BBBP BACE Tox21
NatureLM 1B 0.711 0.794 0.683
NatureLM 8B 0.702 0.820 0.698
[…]
CamS-LLaMA 0.1B 0.942 0.870 0.827
Appendix E Extended Discussion on Data Scale, Fairness, and Efficiency
A prominent distinction between CamS-LLaMA and graph-native baselines (e.g., KPGT) is the pre-training corpus size (675M vs. 550K). We emphasize that this disparity does not constitute experimental unfairness, but rather demonstrates a critical methodological advantage of our sequence-based framework: Scalability.
Graph-native foundation models typically rely on heavy auxiliary inputs or targets to compensate for the lack of semantic density in raw graphs. This creates distinct scaling barriers:
  • Descriptor Calculation (e.g., KPGT): KPGT requires computing over 200 RDKit descriptors and fingerprints for every training instance to serve as ”knowledge” targets. Scaling this dense annotation to the 675M-scale Enamine dataset (let alone augmented views) imposes significant data engineering and storage overheads.
  • 3D Conformer Generation (e.g., GEM, Uni-Mol): Other baselines like GEM explicitly rely on 3D geometric views. Generating conformers (via ETKDG or DFT) allows for richer physics but is computationally prohibitive at the billion-scale.
These pre-processing costs limit such methods to smaller datasets (e.g., $\sim$500K) by design necessity.
In contrast, CamS transforms molecular graphs into causal sequences via purely logical operations (BPE merge and BFS traversal). These operations involve only string manipulation and standard graph traversal, which are CPU-efficient and easily parallelizable. CamS operates directly on pure molecular topology without requiring external descriptor injection or 3D optimization during pre-training. This efficiency removes the data-scaling bottleneck, allowing our method to leverage billion-scale augmented datasets as a standard feature of its training pipeline.
Diversity over Repetition (High-Coverage Training).
Our model was trained for $\sim$1 epoch on the augmented Enamine dataset (5 views per molecule). While baselines typically train on a small dataset (550K) for many epochs (High Repetition), CamS processes a vast number of unique structures (675M) with limited repetition per structure (High Diversity). The performance superiority of CamS, therefore, stems from its ability to efficiently access and learn from the breadth of the chemical space. This ”Diversity over Repetition” regime is unlocked specifically by the efficient nature of the CamS representation, which would be computationally infeasible for descriptor-heavy graph baselines.
Appendix F Detailed Experimental Results
@@
Appendix G Interpretability Details for Activity-Cliff Attention Analysis
Overview. This section provides implementation details for the attention analysis presented in Section 4.3. It complements the main text by specifying: (1) the exact algorithm for identifying differential atoms in activity-cliff pairs (Appendix G.1, Algorithm 7); (2) the mapping from atom-level labels to CamS tokens across different Motif Scales (Appendix G.2); and (3) the detailed computation of the Rel-DTAP metric (Appendix G.3).
G.1 Activity-cliff Pair Construction and Differential/Shared Atom Identification
@@
G.2 Mapping Atom-level Diff/Shared Labels to Tokens in Each Scale Region
@@
Let $\mathcal{A}_{\Delta}$ be the differential-atom set for this molecule from Appendix G.1. We assign a token-level differential indicator by an any-diff rule: a token at position $t$ is marked as differential iff $S_{t}\cap\mathcal{A}_{\Delta}\neq\varnothing$; otherwise it is marked as shared/non-differential. Thus a token covering multiple atoms is counted as differential if it contains at least one differential atom.
@@
G.3 Attention Extraction and Metric Computation
@@
Within a given scale region $s$ (defined by [CONCAT] boundaries), let $p_{i}$ denote the (renormalized) attention weight on token $i$ in that region, and let $d_{i}\in\{0,1\}$ be its differential indicator from Appendix G.2. We define the mean differential-token attention and mean shared-token attention as
@@
G.4 Case-Study Pair Selection Protocol
To ensure the representativeness and reproducibility of the qualitative analysis in Section 4.3, we employ a rigorous, log-driven selection pipeline rather than manual cherry-picking. We build case studies from the test-set activity-cliff pairs identified in Appendix G.1. We implement three complementary selection modes:
  1. 1.
    Mode A: “Similar-structure” cases (Targeting Minimal Edits). This mode isolates pairs that differ by only a few atoms but yield a large potency change.
    • Atom-count prefilter: We require the atom count difference $\Delta N\leq 3$ and pair size $N_{\max}\leq 100$.
    • MCS-based scoring: We compute the MCS and define the edit size $d_{\max}$ as the number of non-MCS atoms. We prioritize pairs with $d_{\max}\leq 5$.
    • Ranking: Pairs are ranked by increasing edit size (more similar first) and then by decreasing fold change.
  2. 2.
    Mode B: “Largest fold-change” cases (Targeting Extreme Cliffs). This mode targets the most extreme potency shifts regardless of edit size. We sort eligible pairs by potency fold-change in descending order and select top candidates.
  3. 3.
    Mode C: “Relatively larger molecules” cases (Targeting Complexity). This mode ensures coverage of complex molecules. We filter for pairs with $N_{\max}\geq 40$ and small $\Delta N$, selecting those with the largest molecule sizes.
Justification of Token-Level Labeling (Holistic Chemical Semantics).
A potential concern is whether the ”any-diff” rule (marking a coarse motif as differential if it contains any modified atom) biases metrics toward larger tokens. We argue this design is grounded in holistic chemical semantics. In medicinal chemistry, modifying a single atom within a ring system or functional group (e.g., H $\to$ F on a benzene ring) fundamentally alters the electronic and steric properties of the entire substructure. In CamS, such a modification results in a completely distinct Token ID for the coarse motif. Therefore, high attention to this ”diff-containing” coarse token reflects a valid recognition of the macro-semantic shift of the functional group as a whole, rather than a metric artifact.
G.5 Note on Baselines.
The Rel-DTAP metric is specifically designed to evaluate the multi-granular attention allocation inherent to the CamS hierarchy (mapping attention to specific $1\mathrm{K}$ vs. $67\mathrm{K}$ token regions). Since standard graph baselines (e.g., KPGT, GROVER) operate on fixed input granularities (atoms or triplets) and lack this explicit hierarchical tokenization, this metric is not directly applicable to them. Thus, our analysis focuses on the intrinsic mechanism of CamS.
G.6 Additional Case Studies
We provide additional visualizations for pairs selected via Mode A (Minimal Edits) and Mode B (Extreme Cliffs) to further substantiate the findings.
Figure 3: Additional Case Study 1 (Minimal Edit). Attention heatmap for pairs (from CHEMBL234Ki) with minimal-atom substitution causing activity cliffs.
Figure 4: Additional Case Study 2 (Extreme Cliff). Attention heatmap for pairs with a scaffold-hopping modification yielding a $4000\times$ shift.
Appendix H Detailed Ablation Analysis
In this section, we provide the complete breakdown of the ablation study.
Budget Matching Protocol. To ensure a fair comparison, all single-scale variants were trained with a computational budget strictly matched to the full model. Specifically, we controlled the total number of parameter update steps (and thus the total number of training samples seen) to be identical across all settings, eliminating training duration as a confounding factor.
We compare the following variants across all tasks:
  • CamS (Full): The proposed multi-scale model with fingerprint injection.
  • w/o FP: The pure multi-scale sequence model without fingerprint injection.
  • 1K Only: Single-scale model using fine-grained motifs.
  • 67K Only: Single-scale model using coarse-grained motifs.
1. Impact of Multi-Scale Concatenation. As shown in Table 11, the CamS (Full) model consistently outperforms the single-scale variants (1K Only and 67K Only). Notably, the 67K Only variant suffers significant performance degradation in regression tasks (AVG RMSE 1.329), confirming that excessive compression reduces supervision density.
2. Pure Sequence vs. Fingerprint Injection. On tasks such as Estrogen and ESOL, the w/o FP variant achieves the best performance (bolded). This indicates that the intrinsic structural features learned by CamS are sometimes superior to explicit fingerprints. However, as shown in Table 12, FP injection is crucial for stability on activity-cliff tasks.
[…]
Task KPGT CamS (Full) w/o FP 1K Only 67K Only
Classification (AUROC $\uparrow$)
BACE 0.855(0.014) 0.870(0.013) 0.850(0.014) 0.845(0.030) 0.837(0.024)
BBBP 0.908(0.012) 0.942(0.015) 0.933(0.021) 0.926(0.016) 0.873(0.013)
ClinTox 0.946(0.026) 0.935(0.017) 0.918(0.004) 0.921(0.021) 0.885(0.023)
Estrogen 0.906(0.034) 0.917(0.050) 0.920(0.046) 0.912(0.056) 0.904(0.069)
Metstab 0.889(0.057) 0.891(0.059) 0.888(0.063) 0.875(0.047) 0.869(0.072)
SIDER 0.649(0.011) 0.655(0.016) 0.647(0.023) 0.646(0.017) 0.654(0.015)
ToxCast 0.745(0.003) 0.724(0.008) 0.723(0.019) 0.717(0.016) 0.706(0.015)
Tox21 0.848(0.017) 0.827(0.028) 0.823(0.032) 0.821(0.024) 0.815(0.016)
AVG (Cls) 0.843 0.845 0.838 0.833 0.818
Regression (RMSE $\downarrow$)
ESOL 0.804(0.102) 0.761(0.046) 0.740(0.054) 0.803(0.045) 0.956(0.045)
FreeSolv 2.121(1.025) 2.110(0.959) 2.192(0.832) 2.188(0.820) 2.328(0.919)
Lipo 0.600(0.012) 0.645(0.023) 0.652(0.022) 0.653(0.027) 0.706(0.014)
AVG (Reg) 1.175 1.172 1.195 1.215 1.329
Table 12: Full Ablation Results on MoleculeACE (RMSE $\downarrow$). The column order is consistent with Table 11. The Full model consistently yields the best stability (lowest RMSE).
Task KPGT CamS (Full) w/o FP 1K Only 67K Only
CHEMBL1862 0.633 0.600 0.635 0.624 0.609
CHEMBL1871 0.605 0.604 0.604 0.606 0.623
CHEMBL2034 0.679 0.619 0.679 0.615 0.655
CHEMBL204 0.666 0.709 0.749 0.730 0.726
CHEMBL2047 0.578 0.519 0.578 0.518 0.563
CHEMBL214 0.652 0.635 0.662 0.660 0.663
CHEMBL2147 0.587 0.577 0.632 0.605 0.596
CHEMBL218 0.625 0.632 0.687 0.655 0.646
CHEMBL219 0.718 0.729 0.736 0.723 0.738
CHEMBL228 0.669 0.669 0.679 0.676 0.713
CHEMBL231 0.610 0.630 0.642 0.642 0.638
CHEMBL233 0.691 0.692 0.719 0.724 0.712
CHEMBL234 0.606 0.624 0.630 0.643 0.645
CHEMBL235 0.624 0.612 0.629 0.604 0.629
CHEMBL236 0.655 0.669 0.709 0.708 0.704
CHEMBL237E 0.716 0.684 0.695 0.712 0.737
CHEMBL237K 0.660 0.659 0.659 0.721 0.712
CHEMBL238 0.537 0.537 0.585 0.564 0.572
CHEMBL239 0.644 0.647 0.672 0.676 0.655
CHEMBL244 0.698 0.696 0.726 0.723 0.725
CHEMBL262 0.627 0.629 0.662 0.637 0.662
CHEMBL264 0.574 0.562 0.570 0.583 0.583
CHEMBL2835 0.373 0.384 0.392 0.385 0.384
CHEMBL287 0.706 0.685 0.685 0.683 0.742
CHEMBL2971 0.571 0.574 0.596 0.584 0.584
CHEMBL3979 0.669 0.639 0.672 0.652 0.655
CHEMBL4005 0.559 0.543 0.581 0.553 0.561
CHEMBL4203 0.820 0.787 0.820 0.800 0.811
CHEMBL4616 0.587 0.538 0.565 0.553 0.565
CHEMBL4792 0.619 0.651 0.659 0.668 0.658
AVG 0.632 0.624 0.650 0.641 0.649